OpenStudy (anonymous):

what is the factor of m^4-y^2-12y-36

OpenStudy (anonymous):

(m^2 - y - 6) (m^2 + y + 6)

OpenStudy (souvik):

=m^4-(y^2+12y+36) =m^4-(y+6)^2 =(m^2+y+6)(m^2-y-6)